### v1

@@ Abstract
keywords:
AI regulation, safety, transparency, non-discrimination, environmental sustainability, traceability
@@ 1 Introduction
In today’s world, where Artificial Intelligence (AI) is swiftly reshaping numerous facets of our daily life, the demand for robust and efficient AI regulations has become more pronounced. Acknowledging this need, on October 30, 2023, U.S. President Joe Biden enacted an Executive Order, signifying a historical move towards tackling the complex issues raised by AI technologies [1]. This action highlights an increasing recognition of the profound influence of AI on aspects such as safety and security, privacy, fairness and civil rights, as well as innovation and competition, calling for a forward-thinking and preemptive strategy in regulatory measures regarding AI advancements.
The evolution of AI regulations represents a contemporary and dynamic narrative that has emerged significantly over recent decades. Initial apprehensions about ethics and data privacy have evolved into more defined guidelines and potential legislative frameworks, though the journey is far from complete. The chronicle of AI regulation is a modern and swiftly developing area, mirroring the rapid advancements in AI technology. The genesis of this regulatory trajectory can be traced back to the mid-20th century, marked notably by the introduction of the Fair Credit Reporting Act in 1970 in the United States [2]. The legislation safeguards data gathered by consumer reporting agencies, including credit bureaus, medical data firms, and tenant screening services. The Act prohibits sharing information in a consumer report with individuals or entities that do not have a designated purpose as outlined in the Act. This act represented an early significant step in the formalization of data protection. Another significant achievement is the General Data Protection Regulation (GDPR) [3], which is a thorough data privacy and security legislation enacted by the European Union (EU) in 2016. It stands as the most stringent data privacy and security law globally. While it was formulated and ratified by the EU, it places responsibilities on organizations worldwide as long as they interact with or gather data from individuals in the EU. The GDPR introduces a set of novel data privacy rights to grant individuals greater authority over the information they provide to organizations. As a visual representation, Figure 1, encapsulates the critical milestones in the progression of data protection regulations.
@@
AI regulation goes beyond the data component to encompass AI systems’ design, development, deployment, and use. It is concerned not only with the data these systems process but also their decisions, the biases they may perpetuate, and the ethical implications of their interaction with humans and other systems. AI regulation aims to address the societal and ethical challenges posed by increasingly autonomous technologies. Regulation of AI varies from country to country, reflecting different approaches, priorities, and concerns. For example, in August 2023, the Australian Human Rights Commission (i.e., Commission) submitted a discussion document titled “Promoting Responsible AI” to the Department of Industry, Science, and Resources in Australia [4]. The Commission expresses particular concern about emerging issues, including privacy breaches, algorithmic discrimination, biases in automation, and the spread of misinformation and disinformation. Furthermore, the intricate interaction between AI, neurotechnology, the metaverse, and extended reality technologies adds complexity to this rapidly evolving field. The Commission has advised the government to initiate an assessment of regulatory deficiencies to ascertain the relevance of existing legislation to AI. In cases where gaps are identified, the associated legislation should be examined and updated to effectively address AI-related challenges.
@@
In the initial stages of research, Scherer [5] initiated a discussion on the feasibility and challenges of government AI regulation, suggesting paths for effective regulation. Concurrently, other researchers delved into AI’s regulatory and ethical aspects in specific domains. For example, Pesapane et al. [6] explored these issues in medical services, while Reddy et al. [7] proposed a governance model for ethical and regulatory concerns in AI healthcare applications. Focusing on particular aspects of AI regulation, Wäschle [8] conducted a systematic literature review on highly automated driving, emphasizing safety assessment methods for AI systems. Other reviews have concentrated on different aspects: reviews [9, 10] on AI transparency; works [11, 12, 13] on bias and discrimination in AI; and reviews [14, 15, 16, 17] on AI’s sustainability.
@@
AI and algorithm-driven decision-making are increasingly impacting our daily lives, being extensively used in critical sectors, such as healthcare, business, government, education, and justice. This shift towards an algorithmic society comes with many benefits and risks, as these systems can sometimes cause harm to users and society. Ensuring the safety, reliability, and trustworthiness of these systems is crucial. Trustworthy AI systems are those that are reliable, ethical, and transparent, enabling users to have confidence in the decisions made by these systems. Recent literature reviews on trustworthy AI [18, 19, 20], including work [20] focused on the elements of fairness, explainability, accountability, reliability, and acceptance to mitigate AI risks and enhance user and societal trust. On the other hand, Li et al. [18] provided specific recommendations for AI practitioners, covering the entire AI system lifecycle, from data collection, model development, and system deployment to ongoing monitoring and governance.
@@
In April 2021, the European Commission proposed the initial AI regulatory framework for the European Union [21]. This framework classifies AI systems into various risk levels, with corresponding levels of regulation. The primary objective of this Act is to guarantee the safety, transparency, traceability, non-discrimination, and environmental sustainability of AI systems used in the EU. The oversight of AI systems should rely on human decision-making (i.e., autonomy) rather than full automation to avoid adverse consequences. Members of the European Parliament endorsed the regulation, which was reached in negotiations with member states in December 2023, with 523 votes in favour, 46 against and 49 abstentions [22].
@@
To the best of our knowledge, there lacks a comprehensive survey that fully addresses the five fundamental principles laid out in the AI Act [21], i.e., safety, transparency, traceability, non-discrimination, and environmental sustainability, as they apply to AI models and systems. An analytical paper is needed to explore how regulations might balance the push for innovation with the need to mitigate risks such as unforeseen consequences or misuse of AI technology. In addition, a detailed review of existing AI regulations would enlighten policymakers, stakeholders, and the public and support more informed decisions in AI governance. Motivated by these considerations, we thoroughly examine AI regulation, exploring the interactions and tensions among its essential aspects. We also provide structured, actionable recommendations for AI professionals in academia and industry. The following is a summary of the main contributions of this paper.
@@ 2.1.1 Ethical and societal reasons
The unbiasedness of the AI system depends on its training data and design decisions. Without proper regulations, these systems risk perpetuating existing societal biases, leading to unfair or discriminatory outcomes. This is particularly concerning in areas such as employment, lending, and criminal justice, where biased AI could reinforce existing inequalities. Regulations are necessary to ensure that AI systems are developed and deployed in a manner that is equitable and fair to all segments of society [23].
The increasing deployment of AI in critical sectors like healthcare, finance, and law enforcement raises important questions about accountability [24]. When an AI system makes a decision that negatively impacts individuals or communities, it is essential to have clear guidelines determining who is responsible – the developers, the users, or the AI itself. Regulations in this area would help clarify lines of responsibility and ensure that affected parties have recourse in the event of harm or injustice.
AI’s reliance on vast datasets, often comprising personal and sensitive information, heightens concerns about data privacy and security [25]. The potential for misuse or unauthorized access to this data poses significant risks to individuals’ privacy rights. Regulations are needed to establish strict guidelines on data collection, use, storage, and sharing. These include ensuring that individuals are informed and consented to how their data is used and implementing robust security measures to protect data from breaches and leaks.
Another reason is to mitigate the risk of social manipulation [26]. In particular, AI’s role in social media algorithms can influence public opinion and even manipulate elections, as evidenced by the Cambridge Analytica scandal, where personal data was used to target voters and potentially affect the political landscape. This has led to calls for tighter regulations on how AI can be used to manipulate information and target users on social media platforms through misinformation and disinformation.
@@ 2.1.2 Technical reasons
In fields such as autonomous driving or medical diagnostics, where AI systems make decisions that can have life-or-death consequences, the importance of risk management cannot be overstated [27]. Regulations are critical in ensuring that these systems are designed, tested, and operated under stringent safety standards. These involve setting benchmarks for performance, establishing protocols for failure detection and response, and ensuring systems are resilient to various types of risks, including cyber threats, system malfunctions, and unexpected environmental conditions. Regulatory frameworks can guide the development and deployment of these systems to prevent accidents and mitigate potential harms.
@@
Furthermore, large AI models typically require substantial computational power, translating into significant energy consumption and consequent environmental impacts [28]. Regulations could play a critical role in pushing for developing more energy-efficient AI technologies, potentially reducing both the environmental footprint and the costs associated with high energy consumption.
@@ 2.1.3 Global and National Security Reasons
AI’s capabilities in processing vast amounts of data make it a powerful tool for surveillance, espionage, and cyber warfare. This raises concerns about the protection of national interests and security [29]. A regulatory framework is necessary to govern the use of AI in these domains, ensuring that it does not compromise national security or infringe on citizens’ rights and freedoms. Regulations can help define permissible uses of AI in surveillance and intelligence, establish safeguards against unauthorized data access, and protect against cyber threats that utilize AI technologies. This regulatory oversight is pivotal in maintaining the delicate balance between the use of AI for national security and the preservation of democratic values and individual rights.
The prospect of AI being used in autonomous weapons systems is a growing concern [30, 31]. These technologies can potentially revolutionize warfare, raising ethical and strategic questions about their deployment. Regulations are crucial in preventing the harmful use of AI in military applications, ensuring that such systems are developed and used in compliance with international humanitarian laws and ethical standards. By establishing clear boundaries and oversight mechanisms, regulations can prevent an arms race in lethal autonomous weapons and ensure that AI is used to enhance, rather than undermine global security and peace.
@@ 2.1.4 Economic Reasons
As AI becomes more integrated into consumer products and services, it is vital to protect consumers from faulty or misleading AI applications. Regulations are crucial in setting standards for AI quality, reliability, and truthfulness in advertising [32]. This includes ensuring that AI-driven products meet certain performance benchmarks and that claims made about AI capabilities are accurate and not misleading. By doing so, regulations safeguard consumers from subpar or potentially harmful AI-driven products and services, fostering trust and confidence in the market.
@@ 2.2.1 Safety
AI safety, a multifaceted and interdisciplinary domain, focuses on preventing adverse consequences that might arise from AI systems [33]. This field tackles not only ethical considerations, ensuring AI systems align with moral values and contribute positively, but also addresses technical challenges [34]. These challenges include monitoring and maintaining the reliability of AI systems to avert risks.
In the EU AI Act, safety refers primarily to the regulation and mitigation of risks associated with using AI systems, ensuring that these systems do not pose unacceptable risks to health, safety, and fundamental rights [35]. Consider the machinery industry as an example. High-risk AI systems here might involve AI-driven robots in manufacturing plants, such as an autonomous robotic arm assembling heavy machinery parts or an AI system managing critical manufacturing processes. The malfunction of these systems can cause severe production setbacks, expensive damages, or even endanger human workers [36]. To mitigate these risks, it’s crucial to equip AI systems with robust safety features like fail-safes or emergency stops. They must undergo comprehensive testing and validation across different operational scenarios. Moreover, resilience against cyber threats or manipulation is vital to prevent potentially disastrous outcomes.
@@ 2.2.2 Transparency
The concept of transparency in AI regulation encompasses the level of openness and comprehensibility in how AI systems are developed, implemented, and managed [29]. For instance, AI systems identified as high-risk must accompany these systems with detailed documentation and instructions that disclose the system’s capabilities, characteristics, and limitations, allowing users to understand and correctly apply the outputs generated by these systems [21].
Beyond these operational aspects, transparency in AI also significantly intersects with the concepts of explainability[37]. This facet of transparency is about demystifying the internal workings of AI systems, providing clarity on how specific decisions are reached. It’s a move towards making AI systems less of a “black box” and more of an open book, where users and stakeholders can understand and rationalize the logic behind AI-driven decisions. This level of transparency is not just a regulatory requirement but a foundation for building trust between users and AI technologies, ensuring that users can interact with these systems confidently and effectively, knowing the rationale behind their outputs and actions [38].
@@ 2.2.3 Non-discrimination
The principle of non-discrimination in the realm of AI is centered on reducing the likelihood of biased decision-making by algorithms [39]. This involves careful attention to the design and the integrity of data sets used in creating AI systems, combined with commitments to rigorous testing, effective risk management, thorough documentation, and consistent human oversight throughout the entire lifecycle of these AI systems.
In the context of AI regulations, non-discrimination is a guiding principle that ensures AI technologies are developed and utilized to prevent biased and unfair outcomes or decisions, particularly those that might unjustly target or disadvantage different groups of people[21]. This aspect of AI regulation is crucial given the risk that AI systems might unintentionally sustain, magnify, or even introduce new forms of biases or discriminatory practices [40]. Such regulations aim to encourage the creation and application of AI technologies in a manner that upholds fairness and ethical standards, thus preventing the reinforcement of existing social disparities. For instance, as outlined in OpenAI’s GPT-4 technical report, content considered harmful encompasses elements like “hate speech, discriminatory language, incitements to violence, or content that is then used to either spread false narratives or to exploit an individual” [41]. In response to this, GPT-4 was developed using a technique known as model-refusal. This approach, grounded in reinforcement learning, rewards the model during its training phase for actively declining to produce such harmful content.

### v2

@@ 1 Introduction
In today’s world, where Artificial Intelligence (AI) is swiftly reshaping numerous facets of our daily life, the demand for robust and efficient AI regulations has become more pronounced. Acknowledging this need, on October 30, 2023, U.S. President Joe Biden enacted an Executive Order, signifying a historical move towards tackling the complex issues raised by AI technologies (?). This action highlights an increasing recognition of the profound influence of AI on aspects such as safety and security, privacy, fairness and civil rights, as well as innovation and competition, calling for a forward-thinking and preemptive strategy in regulatory measures regarding AI advancements.
The evolution of AI regulations represents a contemporary and dynamic narrative that has emerged significantly over recent decades. Initial apprehensions about ethics and data privacy have evolved into more defined guidelines and potential legislative frameworks, though the journey is far from complete. The chronicle of AI regulation is a modern and swiftly developing area, mirroring the rapid advancements in AI technology. The genesis of this regulatory trajectory can be traced back to the mid-20th century, marked notably by the introduction of the Fair Credit Reporting Act in 1970 in the United States (?). The legislation safeguards data gathered by consumer reporting agencies, including credit bureaus, medical data firms, and tenant screening services. The Act prohibits sharing information in a consumer report with individuals or entities that do not have a designated purpose as outlined in the Act. This act represented an early significant step in the formalization of data protection. Another significant achievement is the General Data Protection Regulation (GDPR) (?), which is a thorough data privacy and security legislation enacted by the European Union (EU) in 2016. It stands as the most stringent data privacy and security law globally. While it was formulated and ratified by the EU, it places responsibilities on organizations worldwide as long as they interact with or gather data from individuals in the EU. The GDPR introduces a set of novel data privacy rights to grant individuals greater authority over the information they provide to organizations. As a visual representation, Figure 1, encapsulates the critical milestones in the progression of data protection regulations.
@@
AI regulation goes beyond the data component to encompass AI systems’ design, development, deployment, and use. It is concerned not only with the data these systems process but also their decisions, the biases they may perpetuate, and the ethical implications of their interaction with humans and other systems. AI regulation aims to address the societal and ethical challenges posed by increasingly autonomous technologies. Regulation of AI varies from country to country, reflecting different approaches, priorities, and concerns. For example, in August 2023, the Australian Human Rights Commission (i.e., Commission) submitted a discussion document titled “Promoting Responsible AI” to the Department of Industry, Science, and Resources in Australia (?). The Commission expresses particular concern about emerging issues, including privacy breaches, algorithmic discrimination, biases in automation, and the spread of misinformation and disinformation. Furthermore, the intricate interaction between AI, neurotechnology, the metaverse, and extended reality technologies adds complexity to this rapidly evolving field. The Commission has advised the government to initiate an assessment of regulatory deficiencies to ascertain the relevance of existing legislation to AI. In cases where gaps are identified, the associated legislation should be examined and updated to effectively address AI-related challenges.
@@
In the initial stages of research, Scherer (?) initiated a discussion on the feasibility and challenges of government AI regulation, suggesting paths for effective regulation. Concurrently, other researchers delved into AI’s regulatory and ethical aspects in specific domains. For example, Pesapane et al. (?) explored these issues in medical services, while Reddy et al. (?) proposed a governance model for ethical and regulatory concerns in AI healthcare applications. Focusing on particular aspects of AI regulation, Wäschle (?) conducted a systematic literature review on highly automated driving, emphasizing safety assessment methods for AI systems. Other reviews have concentrated on different aspects: reviews (?, ?) on AI transparency; works (?, ?, ?) on bias and discrimination in AI; and reviews (?, ?, ?, ?) on AI’s sustainability.
@@
AI and algorithm-driven decision-making are increasingly impacting our daily lives, being extensively used in critical sectors, such as healthcare, business, government, education, and justice. This shift towards an algorithmic society comes with many benefits and risks, as these systems can sometimes cause harm to users and society. Ensuring the safety, reliability, and trustworthiness of these systems is crucial. Trustworthy AI systems are those that are reliable, ethical, and transparent, enabling users to have confidence in the decisions made by these systems. Recent literature reviews on trustworthy AI (?, ?, ?), including work (?) focused on the elements of fairness, explainability, accountability, reliability, and acceptance to mitigate AI risks and enhance user and societal trust. On the other hand, Li et al. (?) provided specific recommendations for AI practitioners, covering the entire AI system lifecycle, from data collection, model development, and system deployment to ongoing monitoring and governance.
@@
In April 2021, the European Commission proposed the initial AI regulatory framework for the European Union (?). This framework classifies AI systems into various risk levels, with corresponding levels of regulation. The primary objective of this Act is to guarantee the safety, transparency, traceability, non-discrimination, and environmental sustainability of AI systems used in the EU. The oversight of AI systems should rely on human decision-making (i.e., autonomy) rather than full automation to avoid adverse consequences. Members of the European Parliament endorsed the regulation, which was reached in negotiations with member states in December 2023, with 523 votes in favour, 46 against and 49 abstentions (?).
@@
To the best of our knowledge, there lacks a comprehensive survey that fully addresses the five fundamental principles laid out in the AI Act (?), i.e., safety, transparency, traceability, non-discrimination, and environmental sustainability, as they apply to AI models and systems. An analytical paper is needed to explore how regulations might balance the push for innovation with the need to mitigate risks such as unforeseen consequences or misuse of AI technology. In addition, a detailed review of existing AI regulations would enlighten policymakers, stakeholders, and the public and support more informed decisions in AI governance. Motivated by these considerations, we thoroughly examine AI regulation, exploring the interactions and tensions among its essential aspects. We also provide structured, actionable recommendations for AI professionals in academia and industry. The following is a summary of the main contributions of this paper.
@@ 2.1.1 Ethical and societal reasons
The unbiasedness of the AI system depends on its training data and design decisions. Without proper regulations, these systems risk perpetuating existing societal biases, leading to unfair or discriminatory outcomes. This is particularly concerning in areas such as employment, lending, and criminal justice, where biased AI could reinforce existing inequalities. Regulations are necessary to ensure that AI systems are developed and deployed in a manner that is equitable and fair to all segments of society (?).
The increasing deployment of AI in critical sectors like healthcare, finance, and law enforcement raises important questions about accountability (?). When an AI system makes a decision that negatively impacts individuals or communities, it is essential to have clear guidelines determining who is responsible – the developers, the users, or the AI itself. Regulations in this area would help clarify lines of responsibility and ensure that affected parties have recourse in the event of harm or injustice.
AI’s reliance on vast datasets, often comprising personal and sensitive information, heightens concerns about data privacy and security (?). The potential for misuse or unauthorized access to this data poses significant risks to individuals’ privacy rights. Regulations are needed to establish strict guidelines on data collection, use, storage, and sharing. These include ensuring that individuals are informed and consented to how their data is used and implementing robust security measures to protect data from breaches and leaks.
Another reason is to mitigate the risk of social manipulation (?). In particular, AI’s role in social media algorithms can influence public opinion and even manipulate elections, as evidenced by the Cambridge Analytica scandal, where personal data was used to target voters and potentially affect the political landscape. This has led to calls for tighter regulations on how AI can be used to manipulate information and target users on social media platforms through misinformation and disinformation.
@@ 2.1.2 Technical reasons
In fields such as autonomous driving or medical diagnostics, where AI systems make decisions that can have life-or-death consequences, the importance of risk management cannot be overstated (?). Regulations are critical in ensuring that these systems are designed, tested, and operated under stringent safety standards. These involve setting benchmarks for performance, establishing protocols for failure detection and response, and ensuring systems are resilient to various types of risks, including cyber threats, system malfunctions, and unexpected environmental conditions. Regulatory frameworks can guide the development and deployment of these systems to prevent accidents and mitigate potential harms.
@@
Furthermore, large AI models typically require substantial computational power, translating into significant energy consumption and consequent environmental impacts (?). Regulations could play a critical role in pushing for developing more energy-efficient AI technologies, potentially reducing both the environmental footprint and the costs associated with high energy consumption.
@@ 2.1.3 Global and National Security Reasons
AI’s capabilities in processing vast amounts of data make it a powerful tool for surveillance, espionage, and cyber warfare. This raises concerns about the protection of national interests and security (?). A regulatory framework is necessary to govern the use of AI in these domains, ensuring that it does not compromise national security or infringe on citizens’ rights and freedoms. Regulations can help define permissible uses of AI in surveillance and intelligence, establish safeguards against unauthorized data access, and protect against cyber threats that utilize AI technologies. This regulatory oversight is pivotal in maintaining the delicate balance between the use of AI for national security and the preservation of democratic values and individual rights.
The prospect of AI being used in autonomous weapons systems is a growing concern (?, ?). These technologies can potentially revolutionize warfare, raising ethical and strategic questions about their deployment. Regulations are crucial in preventing the harmful use of AI in military applications, ensuring that such systems are developed and used in compliance with international humanitarian laws and ethical standards. By establishing clear boundaries and oversight mechanisms, regulations can prevent an arms race in lethal autonomous weapons and ensure that AI is used to enhance, rather than undermine global security and peace.
@@ 2.1.4 Economic Reasons
As AI becomes more integrated into consumer products and services, it is vital to protect consumers from faulty or misleading AI applications. Regulations are crucial in setting standards for AI quality, reliability, and truthfulness in advertising (?). This includes ensuring that AI-driven products meet certain performance benchmarks and that claims made about AI capabilities are accurate and not misleading. By doing so, regulations safeguard consumers from subpar or potentially harmful AI-driven products and services, fostering trust and confidence in the market.
@@ 2.2.1 Safety
AI safety, a multifaceted and interdisciplinary domain, focuses on preventing adverse consequences that might arise from AI systems (?). This field tackles not only ethical considerations, ensuring AI systems align with moral values and contribute positively, but also addresses technical challenges (?). These challenges include monitoring and maintaining the reliability of AI systems to avert risks.
In the EU AI Act, safety refers primarily to the regulation and mitigation of risks associated with using AI systems, ensuring that these systems do not pose unacceptable risks to health, safety, and fundamental rights (?). Consider the machinery industry as an example. High-risk AI systems here might involve AI-driven robots in manufacturing plants, such as an autonomous robotic arm assembling heavy machinery parts or an AI system managing critical manufacturing processes. The malfunction of these systems can cause severe production setbacks, expensive damages, or even endanger human workers (?). To mitigate these risks, it’s crucial to equip AI systems with robust safety features like fail-safes or emergency stops. They must undergo comprehensive testing and validation across different operational scenarios. Moreover, resilience against cyber threats or manipulation is vital to prevent potentially disastrous outcomes.
@@ 2.2.2 Transparency
The concept of transparency in AI regulation encompasses the level of openness and comprehensibility in how AI systems are developed, implemented, and managed (?). For instance, AI systems identified as high-risk must accompany these systems with detailed documentation and instructions that disclose the system’s capabilities, characteristics, and limitations, allowing users to understand and correctly apply the outputs generated by these systems (?).
Beyond these operational aspects, transparency in AI also significantly intersects with the concepts of explainability(?). This facet of transparency is about demystifying the internal workings of AI systems, providing clarity on how specific decisions are reached. It’s a move towards making AI systems less of a “black box” and more of an open book, where users and stakeholders can understand and rationalize the logic behind AI-driven decisions. This level of transparency is not just a regulatory requirement but a foundation for building trust between users and AI technologies, ensuring that users can interact with these systems confidently and effectively, knowing the rationale behind their outputs and actions (?).
@@ 2.2.3 Non-discrimination
The principle of non-discrimination in the realm of AI is centered on reducing the likelihood of biased decision-making by algorithms (?). This involves careful attention to the design and the integrity of data sets used in creating AI systems, combined with commitments to rigorous testing, effective risk management, thorough documentation, and consistent human oversight throughout the entire lifecycle of these AI systems.
In the context of AI regulations, non-discrimination is a guiding principle that ensures AI technologies are developed and utilized to prevent biased and unfair outcomes or decisions, particularly those that might unjustly target or disadvantage different groups of people(?). This aspect of AI regulation is crucial given the risk that AI systems might unintentionally sustain, magnify, or even introduce new forms of biases or discriminatory practices (?). Such regulations aim to encourage the creation and application of AI technologies in a manner that upholds fairness and ethical standards, thus preventing the reinforcement of existing social disparities. For instance, as outlined in OpenAI’s GPT-4 technical report, content considered harmful encompasses elements like “hate speech, discriminatory language, incitements to violence, or content that is then used to either spread false narratives or to exploit an individual” (?). In response to this, GPT-4 was developed using a technique known as model-refusal. This approach, grounded in reinforcement learning, rewards the model during its training phase for actively declining to produce such harmful content.